OpenStudy (anonymous):

Please help with this problem.....2x-6 over x-1?? Its a fraction and the answer is 2 I just am lost on how to get there?

OpenStudy (anonymous):

are you trying to solve\[\frac{2x-6}{x-1}=2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well the 2 doesnt show up, my teacher told me thaths the answer. I have to find the common denominator and all that?

OpenStudy (anonymous):

hmmm...then i;m lost. let's start from the beginning.

OpenStudy (anonymous):

what is the original problem?

OpenStudy (anonymous):

ok thank you I will. what sign do you use for a fraction? on here?

OpenStudy (anonymous):

i use the equation editor button...but you can also use parentheses to group the parts of the fraction together

OpenStudy (anonymous):

such as 2x-(6/x-1) would be\[2x-\frac{6}{x-1}\]

OpenStudy (anonymous):

ok so , (2x-6)/(x-1) - (4)/(1-x)

OpenStudy (anonymous):

ahhh..ok. i see now.

OpenStudy (anonymous):

sorry that took a while

OpenStudy (anonymous):

\[\frac{2x-6}{x-1}-\frac{4}{1-x}\]?

OpenStudy (anonymous):

yess yess

OpenStudy (anonymous):

:D ok...here goes

OpenStudy (anonymous):

:) lets see it!

OpenStudy (anonymous):

you understand that x-1 = -1(1-x), right?

OpenStudy (anonymous):

yes I got that

OpenStudy (anonymous):

ok...so i'm going to change the second fraction so that the denominator looks like the first fraction. (get a common denominator)

OpenStudy (anonymous):

ok which is x/1

OpenStudy (anonymous):

\[\frac{2x-6}{x-1}-\frac{4}{1-x}=\frac{2x-6}{x-1}-\frac{4}{-(x-1)}\]

OpenStudy (anonymous):

good so far?

OpenStudy (anonymous):

yes except, when u have to divide by the -1 you only do the denominator?

OpenStudy (anonymous):

it doesn't matter...i can put the negative on top, bottom, or out in front.

OpenStudy (anonymous):

ok got that then, next step is where I get lost

OpenStudy (anonymous):

In fact, I'm going to move it out in front which gives me me minus a negative...which is addition, right?

OpenStudy (anonymous):

so we have\[\frac{2x-6}{x-1}-(-\frac{4}{x-1}) = \frac{2x-6}{x-1}+\frac{4}{x-1}\]

OpenStudy (anonymous):

ok so now its addition, got it

OpenStudy (anonymous):

\[=\frac{2x-6+4}{x-1}\]

OpenStudy (anonymous):

since the denominators are the same...right?

OpenStudy (anonymous):

yes right

OpenStudy (anonymous):

\[=\frac{2x-2}{x-1}=\frac{2(x-1)}{x-1}=2\]

OpenStudy (anonymous):

and as they say in the army..."HOOAH! :)

OpenStudy (anonymous):

yes right

OpenStudy (anonymous):

how do you get the 2(x-1)/(x-1)? Andd yess HOOAH!

OpenStudy (anonymous):

but as we in the Navy say to our army brothers..."Hooah, Shipmate!"

OpenStudy (anonymous):

i factored out a 2 from both terms in the numerator...

OpenStudy (anonymous):

it's distribution in reverse.

OpenStudy (anonymous):

2(x-1) = 2x-2, right?

OpenStudy (anonymous):

ohh ok thank you life saver :) got it! Nicee Navy/army lingo

OpenStudy (anonymous):

my pleasure. :)